OpenStudy (anonymous):

i really suck at these, help ! solve: 2(x+4)-2=4x-2(-3+x)

mathslover (mathslover):

\[\large{\textbf{Hint:}{a(b+c)=ab+ac}}\]

mathslover (mathslover):

what is 2(x+4) @quanna16 ?

OpenStudy (anonymous):

2x+8 @mathslover

OpenStudy (anonymous):

distrib the 2 to get 2x+8-2=4x-2(-3+x) distrib -2 on the right 2x +8-2=4x+6-2x combine like terms 2x+6 = 2x+6

mathslover (mathslover):

good quanna16 now we have : \[\large{2x+8-2=4x-2(-3+x)}\] what is -2(-3+x) ? @quanna16

OpenStudy (anonymous):

6-2x @mathslover

mathslover (mathslover):

good work ... so we have now: \[\large{2x+8-2=4x+6-2x}\] correct?

mathslover (mathslover):

@quanna16 pls reply

OpenStudy (anonymous):

yea @mathslover

mathslover (mathslover):

Ok so can you simplify it?

mathslover (mathslover):

@quanna16 ?

OpenStudy (anonymous):

yeah thanks @mathslover